# Investigating Large Language Models in Diagnosing Students’ Cognitive Skills in Math Problem-solving

### v1

@@ Abstract
[…]
@@ 1 Introduction
Automatic assessment of students’ responses to mathematical problems can support learning at scale. Assessment can serve as feedback to guide students through problem-solving processes, allowing them to refine their reasoning and correct misconceptions as they arise. As teachers’ manual assessment is limited in scale and availability (Stankous, 2016), automated assessment is particularly instrumental in environments where human instructors are unavailable or where student-to-teacher ratios are too high (e.g., MOOCs).
The advancement in AI technologies has expanded the automatic assessment to evaluate more complicated forms of student responses. While early automated assessment methods worked on grading categorical multiple-choice questions, current multimodal and reasoning AI models (e.g., GPT-o1) can process handwritten responses to constructed response questions (Gao et al., 2024; Liu et al., 2024; Chamieh et al., 2024; Baral et al., 2023). As opposed to multiple choices and short answers, these students’ handwritten constructed responses capture students’ thinking processes (Livingston, 2009).
Evaluating students’ constructed responses with AI presents a promising opportunity to diagnose their cognitive skills at scale. Cognitive skills are the abilities to know, apply, and reason mathematical concepts and formulas (Mullis, 2017). Unlike traditional grading that focuses primarily on final answers, cognitive diagnosis enables more granular feedback by identifying where a student’s thinking process may have broken down (Jin et al., 2024; Rahbarnia et al., 2014; Kheong, 1994). For instance, such evaluation can reveal whether a student successfully recalled relevant concepts, represented a problem with an appropriate model, determined a reasonable strategy, and implemented it correctly. These cognitive skills (Mullis, 2017) (complete list in Table 3) often manifest in varied ways through texts and graphics and in subtle ways without overt evidence in students’ responses.
However, such an in-depth diagnosis is challenging because AI models need to understand complicated handwritten responses and need to infer students’ thinking processes beyond visible cues. This diagnosis task involves perceiving graphics in students’ responses, interpreting mathematical notations, and reasoning about students’ problem-solving processes. Existing work has focused on examining AI models’ performance on each of these capabilities in math problem-solving (Didolkar et al., 2024; Fang et al., 2024; Ahn et al., 2024), multimodal perception (Zhang et al., 2024b; a), and reasoning (Wang et al., 2023; Hao et al., 2024) in diverse domains. Whereas previous research has asked how well LLMs can emulate human cognition, our task inverts that perspective, challenging models to infer humans’ logical cognition from partial visual and textual evidence. The diagnosis task is precisely the intersection of these capabilities that makes it particularly complex for current LLMs.
@@
In this work, we investigate how well existing LLMs diagnose students’ cognitive skills in mathematics. Specifically, we address two research questions:
  1. RQ1.
    How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
  2. RQ2.
    How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
@@
To answer these questions, we constructed MathCog, a benchmark dataset designed to evaluate cognitive skill diagnosis. In collaboration with 5 education experts and 15 middle school teachers, we curated 12 middle school math problems, each with 50 student responses diagnosed by teachers based on a diagnostic checklist. Each checklist is based on a well-founded diagnosis framework (Mullis, 2017) and consists of 4 to 6 cognitive skill check items (see Table 1).
[…]
  • Creating a benchmark dataset for evaluating cognitive skill diagnosis performance, MathCog, composed of expert-reviewed diagnostic checklists and teacher-diagnosed results for 639 student responses across 12 middle school math problems.
  • Examining the performance of 16 state-of-the-art closed and open LLMs on diagnosing students’ cognitive skills using MathCog.
  • Investigating the impact of multimodal input, reasoning, and model size on the cognitive skill diagnosis task through a comparison study.
@@ 2 Related Work
We review the foundational framework underlying our cognitive diagnosis task, followed by prior investigations into LLMs’ performance in math-related contexts, and existing benchmarks, highlighting both their relevance and current gaps.
@@ 2.1 Cognitive Diagnosis in Mathematics Education
The TIMSS (Trends in International Mathematics and Science Study) framework (Mullis, 2017) provides a comprehensive and mathematics-specific approach to cognitive diagnosis, comprising content and cognitive domains. The cognitive domain, which evaluates knowledge application, is divided into three key areas: Knowing (recalling definitions, recognizing mathematical entities, computing operations), Applying (determining strategies, representing problems, implementing solutions), and Reasoning (analyzing, integrating, evaluating, justifying). While other frameworks like Bloom’s taxonomy (Bloom et al., 1956) and Polya’s problem-solving framework (Polya, 1957) exist, TIMSS’s mathematics-specific approach, validated through its extensive use across approximately 70 countries since 1995 (Mullis et al., 2023), provides more targeted and reliable guidance for assessing students’ mathematical thinking. Our cognitive skill diagnosis task is founded on this framework that provides deeper insights into students’ mathematical thinking processes, addresses limitations of knowledge-centered assessments, and offers specific guidelines for educational interventions.
@@ 2.2 LLM Capabilities in Mathematical Tasks
LLMs have shown significant progress in mathematical problem-solving, with advanced models achieving improved reasoning capabilities through reinforcement learning (Guo et al., 2025; OpenAI, 2024; Zhong et al., 2024). While larger models and Chain-of-Thought prompting further enhance mathematical reasoning capabilities (Li et al., 2024; Yang et al., 2024a; Wei et al., 2022), current models still struggle with multimodal inputs such as visual elements and handwritten content (Zhang et al., 2024b; Baral et al., 2025; Liu et al., 2024). Although domain-specific models like MathGLM-Vision (Yang et al., 2024b) show promise in addressing these challenges, their performance remains limited by data scarcity (Yan et al., 2024). Beyond direct problem-solving, LLMs need metacognitive abilities to evaluate student responses and identify misconceptions (Baral et al., 2024; Kaggle & Eedi, 2020), yet current research focuses mainly on basic error identification rather than complex cognitive diagnosis. We address a critical gap in the field by providing a comprehensive evaluation of LLMs’ ability to track thinking processes and assess cognitive abilities in mathematical contexts.
2.3 Benchmarks for Evaluating Mathematical Tasks
A diverse array of benchmarks exists for evaluating LLMs’ mathematical capabilities, including GSM8K (Cobbe et al., 2021) for elementary school problems, MATH (Hendrycks et al., 2021) for high school competition questions, MathOdyssey (Fang et al., 2024), and DeepMind’s Mathematics dataset (Saxton et al., 2019). In the educational domain, specialized assessments have emerged, such as MathFish (Lucy et al., 2024), which aligns 9,900 problems with 385 K-12 standards to evaluate models’ capacity to recognize specific skills and concepts, along with handwriting recognition (Baral et al., 2025) and misconception detection in MCQs (Kaggle & Eedi, 2020). However, with few exceptions like (Hellman et al., 2023), there is a notable gap in datasets designed for cognitive diagnosis in open-ended responses. While existing research has mainly focused on simple error identification and scoring, the diagnosis of students’ complex cognitive processes in open-ended questions has been relatively less explored. For our novel cognitive skill diagnosis task, we constructed a dataset that supports a comprehensive evaluation of LLMs’ ability to track implicit thinking processes, identify root causes of errors, and assess students’ cognitive abilities.
@@ 3 Dataset: MathCog
To evaluate LLMs’ performance on the cognitive skill diagnosis, we first created a benchmark dataset composed of secondary school math problems, student responses, diagnostic checklists, and verdicts (Table 1). The math problem and student response data are from AI-Hub [^1] https://www.aihub.or.kr, which contains 5,932 K-12 math problem types and 147,882 handwritten student response images labeled with OCR transcriptions. Each problem type has 5 to 7 isomorphic problems that have the same problem-solving procedures and differ only in the numbers in the problem description (e.g., “Solve $x^{2}+2x-3=0$” and “Solve $x^{2}+x-2=0$”). Among the data, we chose the problem types and student responses corresponding to grades 7-9, as the problems were complex enough to elicit cognitive processes. Among 2,605 problem types from grades 7 to 9, we excluded the problem types that had fewer than 50 student responses to ensure that we had enough student responses to use in our performance evaluation. Finally, we manually reviewed the remaining 35 problem types and removed the problem types whose corresponding isomorphic problems require different mathematical knowledge (e.g., “Solve $x^{2}+2x-3=0$” (quadratic) vs. “Solve $x^{3}+4x-1=0$” (trinomial)). The filtering process left us with 137 problems and 796 student responses from 15 problem types, each involving the same mathematical knowledge and problem-solving procedures. The specific numbers at each stage of filtering are shown in Figure 1.
Figure 1: The dataset creation process and the number of problem types, problems, and student responses at each filtering stage.
@@
When $P=\sqrt{3}$ $-7\sqrt{5}+2\sqrt{3}$, $Q=2\sqrt{3}-\sqrt{5}$ $+2\sqrt{5}$, find the value of $P+Q$. [Uncaptioned image]
@@
realize that if the numbers in
the radicals are the same, he
or she can perform the four
arithmetic operations on the
coefficients with the same
letters?
@@
(e.g., put $\sqrt{5}$ terms together,
etc.) and then find the final
sum?
@@
There is a trapezoid with a lower side length of $8$ $cm$ and a height of $4$ $cm$. If the area of this trapezoid is not less than $28$ $cm^{2}$, find how much more $cm$ the length of the upper side of the trapezoid must be. [Uncaptioned image]
@@
the area of a shape correctly
(e.g., the formula for the area
of a trapezoid)?
@@ 3.1 Diagnostic Checklist
For each problem type, we made a diagnostic checklist commonly applicable to isomorphic problems. These checklists are a list of binary question items mapped to one of the 15 cognitive skills defined in the TIMSS 2019 assessment framework (Mullis, 2017). The authors initially drafted the checklist items and refined them by requesting a review from five math curriculum and evaluation experts with PhD degrees in education and practical experience in making math assessment guidelines (e.g., TIMSS, Korean public school curriculum). Experts gave feedback on the clarity, granularity, and validity of the checklist items. For example, the experts suggested changing “square root term” into “irrational term” for generality and combining items like “Were coefficients calculated correctly?” and “Were addition and subtraction performed correctly?” into “Were the addition and subtraction calculations of coefficients performed accurately?” for reasonable granularity: specific enough to describe a distinct skill, yet general enough to accommodate diverse forms of valid evidence. The experts also commented on the skills each problem can or cannot assess. Experts pointed out that our math problems primarily focus on calculating numbers and applying knowledge and, hence, are limited in assessing “reasoning” (e.g., justifying, analyzing, generalizing) (Mullis, 2017) by their design. We scoped our check items to “knowing” and “applying” cognitive domains only. We took two iterations to refine the checklists, and each checklist was reviewed by two experts independently in each iteration.
@@ 3.2 Teacher-generated Verdict
We recruited 15 middle school math teachers to evaluate 796 student responses based on predefined diagnostic checklists. These teachers had an average of 6.1±4.3 years of experience (range: 2.5–20 years). Each check item was assessed along two dimensions: “yes/no” and “evident/vague” (see Table 1). A “yes” response indicated that the student fully demonstrated the cognitive action specified, while a “no” indicated otherwise. “Evident” meant there was clear evidence to support the judgment, whereas “vague” signified insufficient evidence. For example, for the check item “Did the student calculate correctly?”, teachers marked “evident yes” if the response showed a complete, error-free calculation process, “evident no” if there were clear mistakes, and “vague no” if calculations were missing.
To account for the subjectivity in the diagnosis, we grouped three teachers to evaluate approximately 160 student responses together. The diagnosis occurred in two phases: first, each group collaboratively assessed 36 responses, resolving conflicts through majority voting; second, each teacher independently assessed 18 responses. The second phase included 18 overlapping responses, allowing us to measure inter-rater agreement (see Table  4 in Appendix). Based on the threshold established in prior literature (Graham et al., 2012), we set a minimum agreement of 70% and excluded data from problem types 6, 13, and 14 accordingly. The distribution of the verdicts is presented in Table 5.
@@ 4 Experimental Setting
Using MathCog as a benchmark dataset, we explored two research questions. For RQ1, we evaluated the state-of-the-art LLMs, testing GPT (Achiam et al., 2023), Gemini (Team et al., 2023), Llama (Touvron et al., 2023), Claude (Anthropic, 2024), and DeepSeek (Guo et al., 2025) series. For RQ2, we configured three experiments in which we compared the performance of LLMs in different inputs, reasoning capabilities, and model sizes.
Task Definition. Our task for LLMs is essentially a single-label classification task. Formally, we represent diagnosis result for a student’s response as a collection of tuples $(c,r)$, where $c\in C$ (i.e., check items) and $r\in\{Evident\ Yes,Vague\ Yes,Evident\ No,Vague\ No\}$. The combined judgment $(c,r)$ provides both an evaluative decision and a confidence signal for a skill. Given a context $(P,S,R,C)$, where P is the problem, S is the answer, and R is the student response, the task is to produce assessment tuples $(c,r)$ for all $c\in C$.
@@
Prompting. We instructed LLMs to evaluate each check item based on a given math problem, its solution, a student’s response, and a diagnostic checklist. The LLMs processed inputs in the form of OCR transcriptions, where mathematical formulas and visual cues (e.g., strikethroughs) were represented using LaTeX. In the case of testing the effect of multimodality, we supplied images of student responses. Since the original inputs were in Korean, we machine-translated them into English to prevent possible performance degradation due to language (Achiam et al., 2023). To enhance response accuracy, we employed Chain-of-Thought prompting (Wei et al., 2022; Liu et al., 2024), guiding the LLM to systematically address each check item by first restating its content, identifying relevant evidence, providing an explanation, and delivering a final verdict. The verdict followed one of the four categories used by teachers in MathCog. We opted not to include few-shot examples, as MathCog does not provide explicit evidence and explanations for reference. To minimize randomness, all LLMs were run with a temperature setting of zero. A detailed description of the system and user prompts can be found in the Appendix.
Models. We conducted our experiment using 16 distinct closed and open LLMs from various vendors. To assess the impact of multimodal input, we tested GPT-4o (2024-08-06), Claude-3.5-Sonnet (20240620), and Gemini-1.5-Flash under two conditions: one using text-only prompts and another supplemented with an image. To examine the effect of reasoning capabilities, we compared LLMs with reasoning capabilities—GPT-o1-Preview (2024-09-12), Gemini-2.0-Flash-Thinking (exp-01-21), and DeepSeek-R1—with conventional LLMs of the same model series—GPT-4o, GPT-4o-Mini (2024-07-18), Gemini-1.5-Pro (002), Gemini-1.5-Flash, and DeepSeek-V3 (0324). Finally, we investigated model size differences: GPT-4o, Gemini-1.5-Pro, Llama-3.1-450B (2024-07-23), and DeepSeek-V3 were categorized as large models; Llama-3.3-70B (2024-07-23) and Qwen-2.5-72B (2024-09) as medium models; and GPT-4o-Mini, Gemini-1.5-Flash-8B (001), Llama-3.1-8B (2024-07-23), Qwen-2.5-7B (2024-09), and Mistral-7B (2024-08-22) as small models.
Metrics. We evaluated the outputs of the large language models (LLMs) against ground-truth diagnostic results provided by human experts. Performance was assessed using both macro F1 score and accuracy computed over verdicts (i.e., $(c,r)$) to quantify the effectiveness of skill diagnosis. For a more nuanced analysis, we also examined the models’ tendencies toward overconfidence and underconfidence. Specifically, overconfidence is defined as the conditional proportion of “evident” verdicts produced by the LLMs that were incorrect. Conversely, underconfidence is the conditional proportion of verdicts labeled as “vague” by the LLMs that were, in fact, both correct and should have been classified as “evident.”

### v2

# Benchmarking Large Language Models for Diagnosing Students’ Cognitive Skills from Handwritten Math Work

@@ Abstract
Students’ handwritten math work provides a rich resource for diagnosing cognitive skills, as it captures intermediate reasoning beyond final answers. We investigate how current large language models (LLMs) perform in diagnosing cognitive skills from such work. However, student responses vary widely, often omitting steps or providing only vague, contextually implicit evidence. Despite recent advances in LLMs’ multimodal and reasoning capabilities, their performance under such conditions remains underexplored. To address this gap, we constructed MathCog, a benchmark dataset containing 3,036 diagnostic verdicts across 639 student responses to 110 math problems, annotated by teachers using TIMSS-grounded cognitive skill checklists with evidential strength labels (Evident/Vague). Evaluating 18 LLMs, we find that (1) all models underperform (F1 < 0.5) regardless of capability, and (2) performance degrades sharply under vague evidence. Error analysis reveals systematic patterns: models frequently misattribute vague evidence as evident, overthink minimal cues, and hallucinate nonexistent evidence. We discuss implications for evidence-aware, teacher-in-the-loop designs for LLM-based cognitive diagnosis in educational settings.
@@ 1 Introduction
Diagnosing students’ cognitive skills from their problem-solving work is a long-standing goal in mathematics education, as it can reveal where and how students’ reasoning breaks down beyond final correctness [17, 12]. As defined in the TIMSS assessment framework [22], cognitive skills refer to students’ abilities to know, apply, and reason with mathematical concepts and procedures. Handwritten math work is particularly valuable for diagnosing these skills because it can reflect intermediate reasoning steps and partial understanding [14, 12].
However, such diagnostic evidence is often incomplete, implicit, or unevenly expressed across students [11, 24], making reliable cognitive skill diagnosis inherently challenging. Despite recent advances in large language models (LLMs), including multimodal perception [31, 30] and reasoning capabilities [27, 9], their ability to diagnose students’ cognitive skills from handwritten work remains largely unexamined. While prior work has examined LLMs’ own mathematical problem-solving abilities [5, 6, 2], our work extends beyond problem solving to examine their ability to diagnose human problem-solving processes.
@@
In this work, we systematically investigate how well existing LLMs diagnose students’ cognitive skills in mathematics. Specifically, we focus on the degree of evidential strength in students’ handwritten responses. We address the following research questions:
  • RQ1.
    How do different LLMs (varying in image input, reasoning, model size, and few-shot prompting) perform in diagnosing students’ cognitive skills?
  • RQ2.
    How does the evidential strength of student responses affect LLM’s cognitive skill diagnosis performance?
@@
To answer these questions, we constructed MathCog, an expert-crafted benchmark dataset designed to evaluate cognitive skill diagnosis. In collaboration with 5 education experts and 15 middle school teachers, we curated 12 middle school math topics and 110 problems, each with 50+ student responses diagnosed by teachers based on a problem-specific diagnostic checklist grounded in the TIMSS cognitive framework [22]. Teachers provided binary judgements (Yes/No) for each skill and annotated the presence of supporting evidence as Evident or Vague, yielding 3,036 diagnostic items total. Critically, we curated our dataset to include only problems where human experts achieve >70% agreement, establishing a stable ground truth rather than subjective noise.
Using MathCog, we evaluated 18 closed- and open-source LLMs spanning multiple model families, sizes, and capabilities. Results indicate that current LLMs struggle to reliably diagnose students’ cognitive skills (all F1 scores < 0.5), with performance degrading sharply under vague-evidence conditions and frequently misattributing weak evidence as strong. Qualitative error analysis further reveals the model’s failure patterns across model types. In particular, models often misidentify or hallucinate supporting evidence, or over-infer students’ cognitive skills from incomplete work, showing that errors cascade throughout the diagnostic pipeline from recognizing student responses to interpreting them based on the rubric. Building on these findings, we discuss implications for designing LLM-based diagnostic tools that better elicit students’ diagnostic evidence of cognitive processes and support teacher-in-the-loop interpretation.
Our contributions are threefold. (1) We introduce MathCog, an expert-crafted benchmark dataset for cognitive skill diagnosis. (2) Using MathCog, we evaluate 18 large language models and analyze the effects of multimodality, reasoning, model size, and evidential strength on diagnosis performance. (3) We analyze diagnostic error patterns under vague-evidence conditions and discuss implications for the design and use of LLM-based cognitive diagnostic systems.
@@ 2 Related Work
We review the foundational framework underlying our cognitive diagnosis task, prior investigations into LLMs’ performance in math-related contexts, and existing benchmarks, highlighting their current gaps.
@@ 2.1 Cognitive Diagnosis in Mathematics Education
The TIMSS [22] is a comprehensive and math-specific framework for cognitive diagnosis, comprising content and cognitive domains. The cognitive domain, which evaluates knowledge application, is divided into three key areas: Knowing (recalling definitions, recognizing mathematical entities, classifying or ordering quantities, computing), Applying (determining strategies, representing problems, implementing solution procedures), and Reasoning (analyzing, justifying). This research aims to explore whether LLMs can substitute for the mapping to enable scalable and explainable cognitive diagnosis. To our knowledge, this is the first systematic investigation of LLMs in the context of the TIMSS framework.
@@ 2.2 LLM Capabilities in Mathematical Tasks
Recent advances have enabled LLMs to achieve remarkable performance in mathematical problem-solving [8, 23], yet they show relative weaknesses in diagnosing student abilities [20, 29]. These diagnostic challenges are further complicated by current models’ struggles with multimodal inputs such as visual elements and handwritten content [31, 4, 16]. While some recent work has demonstrated LLMs’ potential for cross-domain cognitive diagnosis [19], existing research has limited systematic exploration of LLMs’ diagnostic capabilities for cognitive skills in math specifically. We address this critical gap in the field by providing the first comprehensive evaluation of LLMs’ ability to diagnose cognitive skills in mathematical contexts using an established educational framework.
2.3 Benchmarks for Evaluating Mathematical Assessment Tasks
In educational settings, several benchmarks have been proposed to evaluate LLMs’ performance in mathematical assessment tasks, including grading, skill recognition, and error analysis. For example, MathFish [18] aligns 9,900 problems with 385 K–12 standards to assess models’ ability to identify targeted mathematical skills and concepts. Automated scoring of open-ended constructed responses in math word problems has also been studied [10]. Other benchmarks focus on misconception detection in multiple-choice questions [13] or handwriting recognition in student work [4]. While prior research has focused on error detection and scoring, diagnosing complex cognitive skills from students’ open-ended handwritten responses remains underexplored and under-resourced. To address this gap, we introduce a novel dataset for cognitive skill diagnosis that enables systematic evaluation of LLMs’ ability to interpret implicit reasoning and diagnose students’ cognitive skills beyond final answers.
@@ 3 Dataset: MathCog
When $P=\sqrt{3}$ $7\sqrt{5}+2\sqrt{3}$, $Q=2\sqrt{3}-\sqrt{5}$ $2\sqrt{5}$, find the value of $P+Q$. [Uncaptioned image]
@@
realize that the coefficients can
be combined when the radicals
are the same?
@@
and then find the final sum?
@@
There is a trapezoid with a lower side length of 88 cmcm and a height of 44 cmcm. If the area of this trapezoid is not less than $28$ $cm^{2}$, find how much more $cm$ the length of the upper side of the trapezoid must be. [Uncaptioned image]
@@
the area of a shape correctly?
@@
To evaluate LLMs’ performance on cognitive skill diagnosis, we created a benchmark dataset comprising secondary-school math problems, handwritten student responses, diagnostic checklists, and teacher-generated verdicts (Table 1). The math problem and student response data are from AI-Hub [^1] The data are available with approval from the provider via AIHub. , which provides OCR transcriptions of handwritten work. Each topic includes multiple isomorphic problems that share the same problem-solving procedures but differ in numerical values (e.g., “Solve $x^{2}+2x-3=0$” and “Solve $x^{2}+x-2=0$”) [25, 21]. From this data, we focused on topics from grades 7–9, where problems are sufficiently complex to elicit students’ cognitive processes. We further excluded topics with fewer than 50 student responses to ensure sufficient data for reliable evaluation. We decided not to augment our dataset with synthesized data, as it might bias our observations on this novel task. Through this filtering process, the resulting dataset contains topics, 137 problems, and 796 student responses.
@@ 3.1 Diagnostic Checklist
For each topic, we developed a diagnostic checklist commonly applicable to isomorphic problems. Each checklist consists of binary question items mapped to one of the 15 cognitive skills defined in the TIMSS 2019 assessment framework [22]. The checklist items were adapted from TIMSS skill descriptions and refined through expert review by five mathematics curriculum and evaluation experts with PhD degrees in education and practical experience in making math assessment guidelines. Experts gave feedback on the clarity, granularity, and validity of the checklist items. The experts also commented on the skills each problem can or cannot assess. Experts pointed out that our math problems primarily focus on calculating numbers and applying knowledge, and hence are limited in assessing “reasoning” (e.g., justifying, analyzing, generalizing) [22] by design. We scoped our check items to “knowing” and “applying” cognitive domains only. We took two iterations to refine the checklists, and each checklist was reviewed by two experts independently in each iteration.
@@ 3.2 Teacher-generated Verdict
We recruited 15 middle school math teachers to evaluate 796 student responses based on predefined diagnostic checklists. Teachers had an average of 6.1±4.3 years of experience (range: 2.5–20 years). Each check item was assessed along two dimensions: correctness (Yes/No) and evidential strength (Evident/Vague). A “Yes” response indicated that the student fully demonstrated the cognitive action specified, while a “No” indicated otherwise, including partially demonstrated responses. “Evident” meant there was clear evidence to support the judgment, whereas “Vague” signified insufficient evidence.
Figure 1: Overview of the diagnostic pipeline of the MathCog.
To account for the subjectivity in cognitive diagnosis, each student’s response was evaluated by three teachers with overlapping assignments, allowing us to measure inter-rater agreement (see Table  7 in Appendix [^2] Appendix can be found: OSF appendix materials). Following the threshold established in prior literature [7], we excluded topics with agreement below 70%, resulting in a final dataset of 12 topics, 110 problems, and 639 student responses. As a result, the benchmark focuses on diagnostic cases with stable expert agreement, providing a reliable basis for evaluating model performance. These topics cover three-fourths of the content domains defined in TIMSS 2019 (see Table 6 in Appendix), and the dataset remains reasonably sized given the substantial human effort required for expert annotation.
@@ 4 Experimental Setting
Using MathCog as a benchmark dataset, we evaluated a diverse set of LLMs with varying input modalities, reasoning capabilities, model sizes, and prompting strategies to address RQ1, and analyzed how the evidential strength of student responses affects diagnosis performance to address RQ2.
@@
Prompting. We instructed LLMs to evaluate each diagnostic check item given a math problem, its solution, a student’s response, and a diagnostic checklist. Student’s response image inputs were provided as OCR transcriptions, with mathematical formulas and visual cues (e.g., strikethroughs) represented in LaTeX. For multimodal settings, we additionally supplied images of student responses. Since the original inputs were in Korean, we machine-translated them into English to prevent possible performance degradation due to language [1]; we used Google Translate API for batch translation and manually verified them. We employed Chain-of-Thought prompting [16] to guide LLMs to systematically address each check item by first restating its content, identifying relevant evidence, providing an explanation, and delivering a final verdict. The verdict followed one of the four categories used by teachers in MathCog. To examine the effect of in-context examples on model performance, we experimented with few-shot prompting using two exemplars that covered all four types of verdicts. To minimize randomness, all LLMs were run with a temperature setting of zero. Full system and user prompts are provided in the Appendix.
Models. We evaluated 18 LLMs from multiple model families, selected to analyze the effects of multimodality, reasoning capability, and model size. For multimodality, selected models were tested under both text-only and image-augmented conditions; for reasoning, we compared reasoning-oriented models with their conventional counterparts within the same model families.
Metrics. We evaluated LLM outputs against teacher-provided ground-truth diagnostic labels. Overall diagnosis performance was assessed using macro F1 score and accuracy over the four verdict categories. We report macro F1 to mitigate the effect of label imbalance across verdicts (see Table 8 in Appendix). To address RQ2, we analyzed how the evidential strength of student responses affects diagnosis performance. To formalize this analysis, we treat each diagnosis instance as a $(\textit{student response},\textit{cognitive skill})$ pair, indexed by $i$. For each instance $i$, the ground-truth labels consist of a skill judgment $y_{i}\in\{\text{Yes},\text{No}\}$ and evidential strength $e_{i}\in\{\text{Evident},\text{Vague}\}$, and the corresponding model predictions $\hat{y}_{i}$ and $\hat{e}_{i}$. To characterize how models handle evidential strength, we define two complementary metrics that capture distinct failure modes in evidence attribution. First, we define evidence over-attribution (OverAttr) as the frequency with which a model assigns “Evident” to cases where the ground-truth evidential strength is “Vague”:
$\text{OverAttr}=\frac{\left|\left\{i\mid e_{i}=\text{Vague}\wedge\hat{e}_{i}=\text{Evident}\right\}\right|}{\left|\left\{i\mid e_{i}=\text{Vague}\right\}\right|}$ (1)
This metric reflects the model’s tendency to overstate evidential support under vague evidence cases, independent of diagnosis correctness. Second, we define evidence false-attribution (FalseAttr) as the frequency with which a model assigns Evident to incorrect diagnoses:
$\text{FalseAttr}=\frac{\left|\left\{i\mid\hat{y}_{i}\neq y_{i}\wedge\hat{e}_{i}=\text{Evident}\right\}\right|}{\left|\left\{i\mid\hat{y}_{i}\neq y_{i}\right\}\right|}$ (2)
This metric captures a particularly misleading failure mode, where erroneous diagnoses are accompanied by strong evidential claims.